OpenStudy (anonymous):

~~~GIVING OUT A MEDAL AND FANNING WHOEVER HELPS THANK YOU!!!~~~ Simplify completely: the quantity 6 times x to the 4th power plus 9 times x to the 2nd power plus 12 times x all over 3 times x 2x3 + 9x2 + 12 2x2 + 3x + 4 2x5 + 3x3 + 4x2 2x3 + 3x + 4

HanAkoSolo (jamierox4ev3r):

6x^4+9x^2+12x/3x

HanAkoSolo (jamierox4ev3r):

then you would use 3x to divide all of the integers in the numerator by the quantity of 3x (ex: 12x/3x=4)

OpenStudy (anonymous):

But i think whats confusing me is the exponents i dont know what to do with them @Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

hint: 9x^2 divided by 3x= 3x

HanAkoSolo (jamierox4ev3r):

based from that, can you figure out the rest?

HanAkoSolo (jamierox4ev3r):

@HelloGoodmorning

OpenStudy (anonymous):

But what about the exponent @Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

6^4/3x=2x^3

OpenStudy (anonymous):

Ok so did you subtract the exponent like 6^4/3^1=2^3

OpenStudy (anonymous):

@Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

well, when you divide, one exponenent is automatically cancelled out

HanAkoSolo (jamierox4ev3r):

so your final answer would be 2x^3+3x+4

OpenStudy (anonymous):

Ohh

OpenStudy (anonymous):

Thank you ^_^

HanAkoSolo (jamierox4ev3r):

no problem ;)